Find the number pattern for 1,3,6,10
1
Expert's answer
2013-02-05T10:00:58-0500
The pattern is a(n) = 1+2+...+n, a(1) = 1, n = 1,2,3...:

n = 1: a(1) = 1 = 1;
n = 2: a(2) = 1+2 = 3;
n = 3: a(3) = 1+2+3 = 6;
n = 4: a(4) = 1+2+3+4 = 10.